# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v4

@@ Appendix G Discussions
G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets—particularly large-scale, user-generated corpora such as LAION-5B [59]—often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
G.2 Toward Linearity of Multimodal Representations

### v5

@@ Appendix I Discussions
I.1 Limitations and Future Directions
Identifying and Quantifying Dataset Biases.
Our study focuses on formalizing the effects of semantic misalignment, but does not directly address the empirical estimation of dataset-specific biases such as selection and perturbation bias. These biases are often latent and difficult to observe, particularly in large-scale, web-curated datasets. Yet, identifying and quantifying them is essential for improving interpretability, data quality, and downstream robustness. This task involves challenges in defining concept taxonomies, determining annotation granularity, and choosing evaluation metrics. Our framework suggests practical strategies: selection bias can be diagnosed via concept coverage statistics over a predefined vocabulary, while perturbation bias may be estimated by comparing dataset captions with those generated by strong image-to-text models, using semantic similarity or factual consistency metrics. We view the development of robust bias identification pipelines, such as those explored by [79], as a valuable direction for future work.
Expanding the Taxonomy of Misalignment.
Multimodal data presents a wide range of potential misalignments across modalities. In this work, we deliberately focus on semantic misalignment in image-text pairs, a particularly prevalent and influential form of bias in vision-language learning. This type of misalignment is central to current multimodal research due to its relevance in large-scale pretraining paradigms and real-world applications (e.g., CLIP, ALIGN). By explicitly modeling selection and perturbation biases, our framework captures common semantic omissions and distortions that significantly affect representational quality. Extending the analysis to other forms of misalignment, such as temporal lag in video-language data, modality-specific dropout, or ambiguity arising from multi-entity co-occurrence, poses additional challenges and may require substantially different modeling assumptions and techniques. We view our proposed abstraction as a tractable and principled foundation for understanding semantic misalignment and believe it can work as a building-block for future explorations on this direction.
Modeling Cross-Modal Semantic Synergy.
Multimodal representations often exhibit semantic synergies arising not only from the alignment of shared content but also from emergent meaning generated through cross-modal integration. At a theoretical level, this aligns with ideas in grounded cognition: cognition (and by extension, semantic representation) is rooted in the integration of perceptual, motor, and introspective modalities via simulation [3]. This parallels our formal distinction between two forms of synergy: i) Some semantics may be deterministically inferable across modalities—such as emotional tone from facial expression, which our latent-space alignment framework readily accommodates. 2) Other higher-order semantics (e.g., sarcasm, irony, humor) emerge through non-additive, nonlinear interactions between modalities and thus require modeling joint-modality latent variables. While our current theory does not explicitly model these emergent semantic phenomena, establishing semantic invariance under structured misalignment is a necessary precursor. Extending our framework to capture structured, emergent cross-modal semantics remains a promising avenue for future research.
Formulating Missing Data in the Latent Space.
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets, particularly large-scale, user-generated corpora such as LAION-5B [61], often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
Linearity of Multimodal Representations.